# The Superconducting Quasiparticle-Amplifying Transmon: A Qubit-Based Sensor for meV Scale Phonons and Single THz Photons

### v3

@@ Abstract
In this letter we propose a novel sensing technology based on weakly charge-coupled transmon qubits, which we call SQUATs: Superconducting Quasiparticle-Amplifying Transmons. The scheme described in this paper differs from the QCD design [6, 7] and that of conventional superconducting qubits [12, 13, 14, 15, 16, 2] by removing the readout resonator entirely from the architecture and relying on direct readout of the qubit transition frequency to detect tunneling events. By removing the readout resonator, we are no longer sensitive to the quantum capacitance as in the QCD, but rather we are sensitive to the quantum inductance in the non-linear LC resonator that makes up the qubit. This architecture change allows for significant reduction in the overall size of the unit cell, increases in pixel density, reduction of two-level system noise, and increase in detection efficiency for uniformly distributed radiation signals. We will show that these devices should outperform the energy sensitivity of competing technologies, while the RF based readout scheme allows them to be naturally multiplexed, thus allowing for highly pixalizable, ultra-low threshold single THz photon and single-phonon detection.
@@
The energy difference between these states, $E_{01}$, is dependent on the qubit’s design, parametrized by the qubit’s single-electron charging energy, $E_{C}=e^{2}/2C$, and the Josephson energy, $E_{J}=\hbar I_{c}/2e$. Here $e$ is the electron charge, $C$ is the qubit capacitance, and $I_{c}$ is the critical current. In the transmon limit, which corresponds to $\xi=\frac{E_{J}}{E_{C}}\gg 1$, we can model the qubit energy accurately as [12]
@@
Figure 1: Schematic of proposed sensor. a) photolithography mask design for prototype sensor. b) QP energy diagram showing signal measurement process. From left to right: A Cooper pair is broken by an incident particle with energy greater than $2\Delta_{\text{Island}}$ and creates QPs in an excited state, these QPs downconvert releasing phonons and lower energy QPs, these QPs diffuse until becoming trapped in the lower gap trap, the QPs then tunnel across the junction until they recombine or are trapped by an impurity. c) Cross-section of the sensor shown on top of a substrate (not to scale). Two possible signal paths are shown: direct absorption of a photon into the island (blue) or an athermal phonon from the substrate (black).
Phonon coupling—Energetic electrons/holes or optical phonons created from an inital particle interaction in the substrate will rapidly downconvert to high energy acoustic athermal phonons. These phonons then undergo anharmonic decay until their mean free path is on the order of the characteristic size of the substrate [17]. The phonons will travel ballistically in the substrate until being absorbed by the active (sensor) and passive (e.g. RF feedline, ground plane) metal films on the surface of the substrate. From [18], the characteristic time scale for these phonon event signals can be approximated by
@@
where $V_{\text{det}}$ is the volume of the substrate, $c_{\text{det}}\right>$ is the average sound speed in the substrate, $A_{i}$ is the area of the $i^{th}$ absorbing material on the detector surface, and $f_{abs}^{i}$ is the phonon transmission probability between the substrate and the absorbing material, modeled simply as
$f_{abs}=1-\exp\left[\frac{2t}{c_{s}\tau_{B}}\right],$ (5)
@@
where $t$ is the absorber film thickness, $c_{s}$ is the sound speed in the film, and $\tau_{B}^{-1}$ is the QP pair breaking rate, values of which can be found in [19]. In our design, the RF feedline and ground plane will be made of Nb, and the majority of the sensing material will be Al.
@@
There are two signal efficiency penalties that must be accounted for in this step. The first is, for a given interaction from an incident particle with an energy greater than the ballistic energy cutoff, the percentage of ballistic phonons with energy greater than twice the superconducting bandgap ($\Delta$) of the sensor material, which is typically $>95\%$ for common materials [20], e.g Si substrate instrumented with Al sensors. Secondly, we must account for the percentage of phonons that are absorbed by the non-instrumented areas of the detector surface. This efficiency factor can be calculated using:
$\varepsilon_{\text{abs}}=\frac{f_{abs}^{\text{active}}A_{\text{active}}}{\sum f_{abs}^{i}A_{i}}.$ (6)
@@
With our initial design we expect to achieve a fill factor of $\sim 5\%$ for instrumented sensor area over the entire chip. The (passive) coplanar waveguide (CPW) feedlines also will cover $\sim 5\%$ of the detector surface. Lastly, a (passive) parquet patterned Nb grid ground plane similar to that used in [21] will cover another $\sim 5\%$ of the surface. Despite the significantly larger $T_{c}$ of Nb vs. Al, preliminary studies suggest that a large fraction of the athermal phonon population in Si will be greater than twice the Nb superconducting gap and thus the Nb ground plane can become a significant source of phonon loss [22]. Using Eq. 6 we expect a phonon collection efficiency due to passive surfaces of $\varepsilon_{\text{abs}}\approx 37\%$. From Eq. 4 the phonon collection time depends linearly on the detector thickness, and for a $500\,\mu\mathrm{m}$ thick Si wafer substrate, we expect $\tau_{\text{phonon}}\approx 2\,\mu\mathrm{s}$. If the signal loss from passive surfaces proves to be a problem, in a future iteration of this design we propose to use a flip-chip process in which the CPW qubit readout and control circuitry is on a separate substrate, similar to what is done in [23]. By doing this, nearly 100% of the passive material can be removed from the detector chip.
@@
Photon coupling— In addition to athermal phonon measurement, these sensors can be used as single photon sensors via direct absorption in the islands (see, e.g., [24, 25]). The shape of the sensor (see Fig. 1) is designed to maximally collect the induced QP signal, as described in the following text; additionally, this shape naturally lends itself to being engineered into a broadband bow-tie antenna.
@@
While a detailed simulation of the antenna structure is beyond the scope of this paper and left for future work, we note that similarly designed antenna micro-structures have achieved broadband absorption in the 1-10 THz range [26, 27, 28]. For example, [27] has shown with simulations $10-100\,\mu\mathrm{m}$-size on-chip bowtie antennas with 20% bandwidth around 1 THz with 90% efficiency and [28] has demonstrated a single-structure bowtie antenna that functions over 1-10 THz. Various studies of THz absorption in general thin films have also been done, with wide-ranging efficiencies [29, 30, 31, 32]. The QCDs achieved an optical efficiency of 90% in a narrow band around 1.5 THz using a mesh antenna absorber [7].
@@
The SQUAT photon absorption efficiency can also be improved by using thicker films, optimizing fabrication methods and material choice, applying surface treatments, and adding a reflector under the substrate to give photons a second pass through the device. From the literature, we conservatively estimate that we can achieve a 50% photon absorption in Al over a 20% bandwidth at THz frequencies [27, 33, 28].
@@
Quasiparticle Collection Efficiency—The coplanar capacitor islands in our device are designed to both maximize surface fill factor and funnel QPs toward the junction. The funneling effect is achieved in part by the choice of a geometry that maximizes the QP collection efficiency for a given sensor area. We can achieve additional collection efficiency through the use of QP trapping [34, 35, 36], similar to that done with Quasiparticle-trap-assisted Electrothermal-feedback Transition-edge sensors (QETs) [37], by fabricating the islands with a material with a larger $T_{c}$ than the junction. For this proposed device, we choose Al for the athermal phonon collection region. For the trapping region and junctions, we require a material with a superconducting gap that is roughly a factor of 10 lower than Al. There are a variety of options for materials that have $T_{c}$’s in the desired range of $\sim 100\,\mathrm{mK}$, and have been shown to grow nice oxide layers for junctions, e.g Hf [38, 39], Ti [40], and AlMn [41, 42, 43, 44]. For the remainder of this concept paper, we will refer to this lower gap material as ‘LG’. The simplest material to use would be AlMn (which we will assume for our calculations here), since its $T_{c}$ in thin films is easily tunable and it lends itself well to Josephson junction fabrication, as it naturally creates an aluminum oxide layer.
@@
    When a phonon or photon of energy greater than $2\Delta_{\text{island}}$ is absorbed by the island, a Cooper pair will be broken, and the resulting QPs will be promoted to well above the bandgap energy. These QPs will undergo a downconversion process to lower energy QPs and phonons, yielding a population of QPs at the band edge after $\mathcal{O}(1)\,\mathrm{ns}$ [45]. During the downconversion process, a portion of the initial event energy will be lost as sub-gap phonons leak into the substrate [19]. For conventional S-wave superconducting films, incident events with energy $E_{\Omega}\gtrsim 4\,\Delta_{\text{island}}$ will have approximately $60\,\%$ of the event energy remain in the QP system [46]. We will assume that the events will be in this energy regime and will adopt a value for the efficiency of the downconversion process ($\varepsilon_{\text{DC}})$ of $60\,\%$, defined as the percent energy in the QP system after downconversion compared to the original event energy.
@@
    The remaining QPs will settle to a stable population with energy of $\sim\Delta_{\text{island}}$ and will diffuse until they either recombine into Cooper pairs over a timescale of $\mathcal{O}(0.1-1)\,\mathrm{ms}$  [47, 48] (in Al), are trapped by impurities in the film, or ideally get trapped in the lower-$T_{c}$ superconductor nearby.
@@
    Once the QPs enter the lower-$T_{c}$ material they undergo a second downconversion process, resulting in an enhancement of QP number that scales with the ratio of the superconducting gaps of the two materials when the ratio is greater than three. These downconverted QPs have an energy lower than $\Delta_{\text{island}}$ and become ‘trapped’. This multiplication process results in a signal enhancement, because the qubit is sensitive to QP number, not energy.
@@
    While the QPs are trapped in the lower $T_{c}$ material, they will tunnel back and forth across the attached Josephson junction until recombination occurs. Each measured tunneling event results in a signal.
A plot of the QP collection fraction ($F_{\text{collect}}$, defined as the measured number of QP tunneling events divided by total number of stable QPs after the down-conversion process) as a function of QP trap and qubit island characteristic lengths, $\ell_{\text{trap}}$ and $\ell_{\text{island}}$ respectively, for a $500\,\mathrm{nm}$ square JJ can be seen in Fig. 2. The details of this calculation are described in Appendix C. Note that while this device adds an additional tunneling step beyond that of its QET counterpart [37], this efficiency loss can be made up by the addition of the QP multiplication and multiple tunnels – allowing for collection fractions that are greater than unity.
Figure 2: Modeled QP collection fraction, $F_{\rm collect}$, as a function of trapping region length and island length for a $500\,\mathrm{nm}$ square Josephson junction with $100\,\mathrm{nm}$ thick Al and LG films. The model accounts for multiple QP tunneling events across the junction, which in principle can result in an efficiency greater than unity.
@@
We model the qubits as notch filters, assuming an ideal transmision line, as [49],
@@
where $f_{bw}$ is the readout bandwidth (inverse integration time), $T_{n}$ is the system noise temperature, and $f_{0}$ is the readout frequency of the SQUAT. Readout efficiency, $\epsilon_{r}$, and quantum noise corrections, $\eta$, are described in Appendix B. For $T_{n}\sim$1.8 K and $f_{0}\sim$6 GHz, Eq. 14 dictates that we can achieve high readout fidelity above -135 dBm readout power for an integration time corresponding to $f_{bw}=10^{4}\,\mathrm{Hz}$. With a quantum-limited readout, we find the reduced equation
@@
Figure 3: Simulation of the signal readout with $\chi=10\,\mathrm{MHz}$ and an effective quality factor of $Q=1\,\mathrm{k}$. Left) Magnitude of $S_{21}$ showing the frequency separation between the odd and even parity states. Right) Charge dispersion of the sensor normalized to the readout frequency, showing the even and odd states. Shown in grey is the target dispersion ($\chi_{0}$). Bottom) Noisy simulation of single QP tunneling background events with an example dark count rate of 1 Hz.
@@
Estimated Sensor Performance—We can combine all of our known efficiency factors to predict the final energy sensitivity of the device to both direct absorption of photons and measurement of athermal phonons in the substrate. Recall that we are measuring total QP number $n_{\text{qp}}\approx\eta_{\text{qp}}/\Delta_{\text{island}}$, where $\eta_{\text{qp}}$ is the total energy in the QP system. The energy in the QP system can be written as the product of the impinging particle energy $E_{\text{particle}}$ (either from a photon or phonon), the absorption efficiency ($\varepsilon_{\text{abs}}$), the efficiency from the downconversion process ($\varepsilon_{\text{DC}}$), and remaining efficiency factors described below. Thus, our measured number of QP tunneling events ($N_{\text{QP}}$) will be
@@
$N_{\text{QP}}=E_{\text{particle}}\times\varepsilon_{\text{abs}}\times\varepsilon_{\text{DC}}\times\frac{1}{\Delta_{\text{Al}}}\times F_{\text{collect}}\times\varepsilon_{\text{BW}}\times\mathcal{F},$ (16)
@@
where $F_{\text{collect}}$ is the QP collection fraction given by Eq. 60, $\varepsilon_{\text{abs}}$ is the efficiency of the particle absorption in the sensor, i.e. the antenna coupling for photons or the signal loss due to passive surfaces for the phonon case, $\varepsilon_{\text{BW}}$ is the measurement efficiency due to the finite sensor bandwidth from Eq. 66, and $\mathcal{F}$ is the readout fidelity given by Eq. 48. Considering only the low energy threshold and operating at with a readout bandwidth of $\sim 10\,\mathrm{kHz}$ with standard RF readout hardware, we can ignore dynamic range effects from the finite bandwidth and set $\varepsilon_{\text{BW}}=1$ and $\mathcal{F}=1$ (see Appendix D and Appendix B for more details on the effects of bandwidth on the dynamic range and readout fidelity of the SQUAT.)
@@
Considering first the athermal phonon detection, taking the worst case scenario that all the phonons are above the Nb gap, our phonon collection efficiency due to passive surface losses is roughly $\varepsilon_{\text{abs}}\approx 37\,\%$. Note that this value is relatively independent of individual sensor size, and the total sensor number can be changed to keep the surface coverage the same. Assuming a readout mode such that single QPs can be counted, Eq. 16 can be converted into an effective phonon energy resolution by setting $N_{\text{QP}}=1$ and solving for $E_{\text{particle}}$. This can then be solved for a variety of sensor geometry parameters by using Eq. 6, Eq. 60, and $\varepsilon_{\text{DC}}=60\,\%$, with the parameters listed in Appendix C.
@@
In Fig. 4 we show the the expected athermal phonon energy sensitivity threshold as a function of sensor island and trap length (see Fig. 1 for definitions) instrumented on a $500\,\mu\mathrm{m}$ thick Si wafer. From the figure, we can see that for various geometries, the SQUAT can easily achieve single meV phonon thresholds. This represents the underlying fundamental sensitivity for the SQUAT, but is in reality the lower bound of detection threshold. In practice, the threshold will ultimately be limited by the dark rate from excess non-equilibrium QPs tunneling across the junction. While dark rates will always be a limitation for counting based sensors, transmon qubits have achieved dark rates as low as $\mathcal{O}(10-100)\,\mathrm{mHz}$ [51], sufficiently low enough to enable us to operate in the single to few QP counting regime.

### v4

@@ Abstract
In this letter we propose a novel sensing technology based on weakly charge-coupled transmon qubits, which we call SQUATs: Superconducting Quasiparticle-Amplifying Transmons. The scheme described in this paper differs from the QCD design [6, 7] and that of conventional superconducting qubits [12, 13, 14, 15, 16, 2] by removing the readout resonator entirely from the architecture and relying on direct readout of the qubit transition frequency to detect tunneling events. By removing the readout resonator, we are no longer sensitive to the quantum capacitance as in the QCD, but rather we are sensitive to the quantum inductance in the non-linear LC resonator that makes up the qubit. This architecture change allows for significant reduction in the overall size of the unit cell, increases in pixel density, reduction of two-level system noise, and increase in detection efficiency for uniformly distributed radiation signals. We will show that these devices should outperform the energy sensitivity of competing technologies, while the RF based readout scheme allows them to be naturally multiplexed, thus allowing for highly pixelizable, ultra-low threshold single THz photon and single-phonon detection.
@@
The energy difference between these states, $E_{01}$, is dependent on the qubit’s design, parametrized by the qubit’s single-electron charging energy, $E_{C}=e^{2}/2C$, and the Josephson energy, $E_{J}=\hbar I_{c}/2e$. Here $e$ is the electron charge, $C$ is the qubit capacitance, and $I_{c}$ is the critical current. In the transmon limit, which corresponds to $\xi\equiv\frac{E_{J}}{E_{C}}\gg 1$, we can model the qubit energy accurately as [12]
@@
Figure 1: a) Photolithographic mask design for a first prototype SQUAT. The islands are composed of two sectors of a circle with island length ($\ell_{\text{island}}$) defined as the radius of the circle. The trap is the region of overlap between the island and the junction material. The length of the overlap is designed to be the same as the width such that the area of the trap can be parameterized with a single length parameter ($\ell_{\text{trap}}$). b) QP energy diagram showing signal measurement process. From left to right: A Cooper pair is broken by an incident particle with energy greater than $2\Delta_{\text{Island}}$ and creates QPs in an excited state, these QPs downconvert releasing phonons and lower energy QPs, these QPs diffuse until becoming trapped in the lower gap trap, the QPs then tunnel across the junction until they recombine or are trapped by an impurity. c) Cross-section of the sensor shown on top of a substrate (not to scale). Two possible signal paths are shown: direct absorption of a photon into the island (blue) or an athermal phonon from the substrate (black).
Phonon coupling—Energetic electrons/holes or optical phonons created from an inital particle interaction in the substrate will rapidly downconvert to high energy acoustic athermal phonons. These phonons then undergo anharmonic decay until their mean free path is on the order of the characteristic size of the substrate [17]. The phonons will then travel ballistically in the substrate until being absorbed by the active (sensor) and passive (e.g. RF feedline, ground plane) metal films on the surface of the substrate. From [18], the characteristic time scale for these phonon event signals can be approximated by
@@
where $V_{\text{det}}$ is the volume of the substrate, $c_{\text{det}}\right>$ is the average sound speed in the substrate, $A_{i}$ is the area of the $i^{th}$ absorbing material on the detector surface, and $f_{abs}^{i}$ is the phonon transmission probability between the substrate and the absorbing material. This transmission probability is modeled simply as
$f_{abs}^{i}=1-\exp\left[\frac{2h^{i}}{c_{s}^{i}\tau_{B}^{i}}\right],$ (5)
@@
where $h^{i}$ is the absorber film thickness, $c_{s}^{i}$ is the sound speed in the film, and $1/\tau_{B}^{i}$ is the QP pair breaking rate, all for the $i^{th}$ film. In our design, the RF feedline and ground plane will be made of Nb, and the majority of the sensing material will be Al, values of $1/\tau_{B}^{i}$ for both of these materials can be found in [19].
@@
There are two signal efficiency penalties that must be accounted for in this step. The first is, for a given interaction from an incident particle with an energy greater than the ballistic energy cutoff, the percentage of ballistic phonons with energy greater than twice the superconducting bandgap ($\Delta$) of the sensor material, which is typically $>98\%$ for common materials [20], e.g Si substrate instrumented with Al sensors. For this reason, we do not consider this efficiency at this point in our simulations. Secondly, we must account for the percentage of phonons that are absorbed by the non-instrumented areas of the detector surface. This efficiency factor can be calculated as the sum of the active (sensor) phonon absorption probability and area, divided by the total phonon absorption probability and area of every phonon absorbing surface:
$\varepsilon_{\text{abs}}=\frac{\sum^{\text{active}}f_{abs}^{i}A_{i}}{\sum^{\text{total}}f_{abs}^{i}A_{i}},$ (6)
@@
From the unit cell for our sensor shown in Fig. 1, we expect this can be placed in an array in such a way that $\sim 5\%$ of the surface area of the entire chip will be active instrumented sensor area. The (passive) coplanar waveguide (CPW) feedlines also will cover an additional $\sim 5\%$ of the detector surface. Lastly, a (passive) parquet patterned Nb grid ground plane similar to that used in [21] will cover another $\sim 5\%$ of the surface. Despite the significantly larger superconducting critical temperature ($T_{c}$) of Nb vs. Al, preliminary studies suggest that a large fraction of the athermal phonon population in Si will be greater than twice the Nb superconducting gap and thus the Nb ground plane can become a significant source of phonon loss [22].
The remaining $\sim 85\%$ of the surface area will be left bare. Various surface treatments (e.g. polishing, chemical etching, etc) will be used in the fabrication R&D to minimize phonon losses at the surfaces. Using these fill factors for the passive and active areas, and weighting by the phonon transmission probabilities in Eq. 5, we thus find from Eq. 6 that we expect a phonon collection efficiency of $\varepsilon_{\text{abs}}\approx 37\%$. From Eq. 4 the phonon collection time depends linearly on the detector thickness, and for a $500\,\mu\mathrm{m}$ thick Si wafer substrate, we expect $\tau_{\text{phonon}}\approx 2\,\mu\mathrm{s}$. If the signal loss from passive surfaces proves to be a problem, in a future iteration of this design we propose to use a flip-chip process in which the CPW qubit readout and control circuitry is on a separate substrate, similar to what is done in [23]. By doing this, nearly 100% of the passive material can be removed from the detector chip.
@@
Photon coupling— In addition to athermal phonon measurement, these sensors can be used as single photon sensors via direct absorption in the islands or coupling of antenna mode to impedance-matched junctions (see, e.g., [24, 25]). The shape of the sensor (see Fig. 1) is designed to maximally collect the induced QP signal, as described in the following text; additionally, this shape naturally lends itself to being engineered into a broadband bow-tie antenna. Analogous to the phonon collection case, $\varepsilon_{\text{abs}}$ is now defined as the internal quantum efficiency for photon absorption, and the characteristic signal is a delta function impulse of energy into the sensor, rather than the exponential signal with a time constant defined in Eq. 4.
@@
While a detailed simulation of the antenna structure is beyond the scope of this paper and left for future work, we note that similarly designed antenna micro-structures have achieved broadband absorption in the 1-10 THz range [26, 27, 28]. For example, [27] has shown with simulations $10-100\,\mu\mathrm{m}$-size on-chip bowtie antennas with 20% bandwidth around 1 THz with 90% efficiency and [28] has demonstrated a single-structure bowtie antenna that functions over 1-10 THz. Various studies of THz absorption in general thin films have also been done, with wide-ranging efficiencies [29, 30, 31, 32]. QCDs have achieved an optical efficiency of 90% in a narrow band around 1.5 THz using a mesh antenna absorber [7].
@@
The SQUAT photon absorption efficiency can also be improved by using thicker films, optimizing fabrication methods and material choice, applying surface treatments, and adding a reflector under the substrate to give photons a second pass through the device. From the literature, we conservatively estimate that we can achieve a photon absorption efficiency of $\varepsilon_{\text{abs}}\approx 50\%$ photon absorption in Al over a 20% bandwidth at THz frequencies [27, 33, 28].
Quasiparticle Collection Efficiency— To maximize the sensitivity of the SQUAT, we must optimize its efficiency to collect QPs. For the SQUAT signal path, ‘collecting a QP’ means that it tunnels across the junction. A novel feature of the SQUAT design is that it acts as a QP multiplier by increasing the number of QPs through downconversion in the trap and through each QP having multiple tunnels, which is further increased by being trapped near the junction. These compounding effects greatly enhance the signal.
The geometry of the coplanar capacitor islands in our device is designed to both maximize surface fill factor and QP collection efficiency. This is optimized in three ways. 1) The geometry of the islands is set by the QP diffusion length of the material such that QPs have a long enough lifetime to reach the junctions. 2) We add QP traps around the junction to provide additional collection efficiency. QP trapping [34, 35, 36], similar to that done with Quasiparticle-trap-assisted Electrothermal-feedback Transition-edge sensors (QETs) [37], can be achieved by fabricating the islands with a material with a larger $T_{c}$ than the junction. 3) With a sufficient difference in the $T_{c}$’s of the island and junction material, the trapping effect also results in a QP multiplication stage, as originally proposed by Booth for superconducting tunnel junctions (STJs) [34].
@@
For this proposed device, we choose Al for the athermal phonon collection region. For the trapping region and junctions, we target a material with a superconducting gap that is roughly a factor of 10 lower than Al. There are a variety of options for materials that have $T_{c}$’s in the desired range of $\sim 100\,\mathrm{mK}$ and have been shown to grow nice oxide layers for junctions, e.g Hf [38, 39], Ti [40], and AlMn [41, 42, 43, 44]. For the remainder of this concept paper, we will refer to this lower gap material as ‘LG’. The simplest material to use would be AlMn (which we will assume for our calculations here), since its $T_{c}$ in thin films is easily tunable by changing the Mn content and it lends itself well to Josephson junction fabrication, as it naturally creates an aluminum oxide layer.
@@
    Phonon collection and initial QP downconversion in island– When a phonon or photon of energy greater than $2\Delta_{\text{island}}$ enters the island, a Cooper pair will be broken with a probability given by Eq. 5, and the resulting QPs will be promoted to well above the bandgap energy. These QPs will undergo a downconversion process to lower energy QPs and phonons, eventually yielding a semi-stable population of QPs at the band edge after $\mathcal{O}(1)\,\mathrm{ns}$ [45]. During the downconversion process, a portion of the initial event energy will be lost as sub-gap phonons leak into the substrate [19]. For conventional S-wave superconducting films, incident events with energy $E_{\Omega}\gtrsim 4\,\Delta_{\text{island}}$ will have approximately $60\,\%$ of the event energy remain in the QP system [46]. We will assume that the initial primary phonon/photon events will be in this energy regime and will adopt a value for the efficiency of the downconversion process ($\varepsilon_{\text{DC}})$ of $60\,\%$, defined as the percent energy in the QP system after downconversion compared to the original event energy.
@@
    QP trapping– The remaining QPs will settle to a stable population with energy of $\sim\Delta_{\text{island}}$ and will diffuse until they either recombine into Cooper pairs over a timescale of $\mathcal{O}(0.1-1)\,\mathrm{ms}$  [47, 48] (in Al), are trapped by impurities in the film, or ideally get trapped in the lower-$T_{c}$ superconductor nearby. A QP has an opportunity to become trapped once it diffuses to the overlapping region between the island the LG material, called the ‘trap’. When the QP enters the trap, it has an energy of roughly $\Delta_{\text{island}}$, which is now much larger than the effective gap of the trap at $\Delta_{\text{trap}}$. These ‘high energy’ QPs now have more available energy states to decay into and a second downconversion process takes place. Once this downconversion process begins in the trap, the average QP energy is now less than $\Delta_{\text{island}}$ and the QPs can no longer enter the island—effectively confining them to the volume of the trap.
@@
    QP multiplication– In addition to the trapping effect described above, we get an additional benefit from the QP downconversion of the enhancement in QP number. When the high energy QPs downconvert into phonons and lower energy QPs, there is a net energy loss due to a percentage of sub-gap phonons just as was the case in the islands, however as long as the ratio of the two superconducting gaps is greater than three, there is a net increase in the total QP number. Since the SQUAT is sensitive to total QP number, this results in a net signal increase. This effect is illustrated by a simple Monte Carlo simulation shown in Fig. 2 (the details of which are described in Appendix C), showing that for a suitable island to trap SC gap ratio, a gain of 5-10 is achievable.
@@
    Multiple measurements of QPs– Finally, we have an additional multiplication step resulting from the ability to measure the same QP multiple times. Once in the trap, the QPs are confined to a small volume near the Josephson junction. If the tunneling time for the particles is much less than the QP lifetime, the QPs will tunnel back and forth across the junction multiple times until recombination occurs. Each tunneling event results in a measurable signal.
To quantify the efficiency of the above process, we define the QP collection fraction as our metric to optimize our design. The collection fraction, $F_{\text{collect}}$, is defined as the ratio of the measured number of QP tunneling events across the Josephson junction to the total number of stable QPs remaining after the first downconversion process in the island, i.e. $F_{\text{collect}}$ quantifies the efficiencies of steps 2-4 above. In defining the collection fraction in this way, we are removing the energy loss from the initial downconversion process ($\varepsilon_{\text{DC}}$) since this parameter is independent of the design itself. For details on exactly how the modeling for $F_{\text{collect}}$ is done, the reader is referred to Appendix C.
A plot of the QP collection fraction as a function of QP trap and qubit island characteristic lengths, $\ell_{\text{trap}}$ and $\ell_{\text{island}}$ respectively, for a $500\,\mathrm{nm}$ square JJ can be seen in Fig. 2. This figure highlights the utility of this sensing mechanism; despite the fact that this device adds the complication of the QP tunneling step beyond that of its QET counterpart [37], this efficiency loss can be made up by the addition of the QP multiplication and measurement of multiple tunnel events – allowing for collection fractions that are greater than unity.
Figure 2: Top: Left Axis– Simulated percentage of energy remaining in the QP system ($\eta_{\text{QP}}$) as a function of incident particle energy relative to the trap gap. In this case the incident particle is a QP of energy $\Delta_{\text{Island}}$, in units of the trapping material’s gap $\Delta_{\text{Trap}}$. Right Axis– Simulated QP number gain, Eq. 61, as a function of SC gap ratio using $\eta_{\text{QP}}$ from the same figure. Bottom: Modeled QP collection fraction, $F_{\rm collect}$, as a function of trapping region length and island length for a $500\,\mathrm{nm}$ square Josephson junction with $100\,\mathrm{nm}$ thick Al and LG films. The model accounts for multiple QP tunneling events across the junction, which in principle can result in an efficiency greater than unity.
@@
We model the qubits as notch filters, assuming an ideal transmission line, as [49],
@@
where $f_{bw}$ is the readout bandwidth (inverse integration time), $T_{n}$ is the system noise temperature, $f_{0}$ is the readout frequency of the SQUAT, and $\sigma^{2}$ is the variance. Readout efficiency, $\epsilon_{r}$, and quantum noise corrections, $\eta$, are described in Appendix B. For $T_{n}\sim$1.8 K and $f_{0}\sim$6 GHz, Eq. 14 dictates that we can achieve high readout fidelity above -135 dBm readout power for an integration time corresponding to $f_{bw}=10^{4}\,\mathrm{Hz}$. With a quantum-limited readout, we find the reduced equation
@@
Figure 3: Simulation of the signal readout with $\chi_{0}=10\,\mathrm{MHz}$ and an effective quality factor of $Q=1000$. Left) Magnitude of $S_{21}$ showing the frequency separation between the odd and even parity states. Right) Charge dispersion of the sensor normalized to the readout frequency, showing the even and odd states. Shown in grey is the target dispersion ($\chi_{0}$). Bottom) Noisy simulation of single QP tunneling background events with an example dark count rate of 1 Hz.
Figure 4: Schematic flow chart showing all the efficiency and conversion factors going from incident particle energy ($E_{\text{particle}}$) to total number of measured QPs ($N_{\text{QP}}^{\text{total}}$). Each efficiency factor is defined as fraction of the measured quantity that is output by the process in consideration, divided by the amount that is left over from the previous process, e.g. output vs input from left to right. In this way, all factors can be multiplied. The definitions for each efficiency can be found: $\varepsilon_{\text{abs}}$: Eq. 6 (for the phonon case), $\varepsilon_{\text{DC}}=60\%$, $F_{\text{collect}}$: Eq. 69 (with definitions there within), $\varepsilon_{\text{BW}}$: Eq. 75 phonon (Eq. 74) (photon), and $\mathcal{F}$: Eq. 48.
@@
Estimated Sensor Performance—We can combine all of our known efficiency factors to predict the final energy sensitivity of the device to both direct absorption of photons and measurement of athermal phonons in the substrate. Recall that we are measuring total QP number $n_{\text{qp}}\approx\eta_{\text{qp}}/\Delta_{\text{island}}$, where $\eta_{\text{qp}}$ is the total energy in the QP system. The energy in the QP system can be written as the product of the impinging particle energy $E_{\text{particle}}$ (either from a photon or phonon), the absorption efficiency ($\varepsilon_{\text{abs}}$), the efficiency from the downconversion process ($\varepsilon_{\text{DC}}$), and remaining efficiency factors described below. Thus, our measured number of QP tunneling events ($N_{\text{QP}}^{\text{total}}$) will be
@@
$N_{\text{QP}}^{\text{total}}=E_{\text{particle}}\times\varepsilon_{\text{abs}}\times\varepsilon_{\text{DC}}\times\frac{1}{\Delta_{\text{Island}}}\times F_{\text{collect}}\times\varepsilon_{\text{BW}}\times\mathcal{F},$ (16)
@@
where $F_{\text{collect}}$ is the QP collection fraction given by Eq. 69, $\varepsilon_{\text{abs}}$ is the efficiency of the particle absorption in the sensor, i.e. the antenna coupling for photons or the signal loss due to passive surfaces for the phonon case, $\varepsilon_{\text{BW}}$ is the measurement efficiency due to the finite sensor bandwidth from Eq. 75, and $\mathcal{F}$ is the readout fidelity given by Eq. 48 (See also Fig. 4 for a pictorial flow of how the output signal is calculated). Considering only the low energy threshold and operating with a readout bandwidth of $\sim 10\,\mathrm{kHz}$ with standard RF readout hardware, we can ignore dynamic range effects from the finite bandwidth and set $\varepsilon_{\text{BW}}=1$ and $\mathcal{F}=1$ (see Appendix D and Appendix B for more details on the effects of bandwidth on the dynamic range and readout fidelity of the SQUAT.)
@@
Considering first the athermal phonon detection, taking the worst case scenario that all the phonons are above the Nb gap, our phonon collection efficiency due to passive surface losses is roughly $\varepsilon_{\text{abs}}\approx 37\,\%$. Note that this value is relatively independent of individual sensor size, and the total sensor number can be changed to keep the surface coverage the same. Assuming a readout mode such that single QPs can be counted, Eq. 16 can be converted into an effective phonon energy resolution by setting $N_{\text{QP}}^{\text{total}}=1$ and solving for $E_{\text{particle}}$. This can then be solved for a variety of sensor geometry parameters by using Eq. 6, Eq. 69, and $\varepsilon_{\text{DC}}=60\,\%$, with the parameters listed in Appendix C.
@@
In Fig. 5 we show the expected athermal phonon energy sensitivity threshold as a function of sensor island and trap length (see Fig. 1 for definitions) instrumented on a $500\,\mu\mathrm{m}$ thick Si wafer. From the figure, we can see that for various geometries, the SQUAT can easily achieve single meV phonon thresholds. This represents the underlying fundamental sensitivity for the SQUAT, but is in reality the lower bound of detection threshold. In practice, the threshold will ultimately be limited by the dark rate from excess non-equilibrium QPs tunneling across the junction. While dark rates will always be a limitation for counting based sensors, transmon qubits have achieved dark rates as low as $\mathcal{O}(10-100)\,\mathrm{mHz}$ [51], sufficiently low enough to enable us to operate in the single to few QP counting regime.